Can you help me solve this math problem: A man can drive a motorboat 45 miles down the Rock River in the same amount of time that he can drive it 27 miles upstream. Find the speed of the current if the speed of the boat is 12 mph in still water.
Answer by scott8148(6628)

let c=current __ t=d/r

45/(12+c)=27/(12-c) __ multiplying by (c+12)(c-12) __ 540-45c=324+27c __ 216=72c __ 3=c